# Normalization in Attention Dynamics

### v1

@@ Appendix A Symmetric initialization
Post-LN. The ODE governing the evolution of the cosine similarity $\gamma(t)$ was already derived in (Geshkovski et al.,, 2025, Theorem 6.8). It is given by
@@
LN-Scaling. The ODE governing $\gamma(t)$ is given by
@@
  • for LN-Scaling $s_{k}\equiv\sqrt{t}$.
@@
  • $\mathcal{V}=e^{-4\sqrt{t}(1+o(1))}$ for LN-Scaling.
@@ Appendix D Final convergence
In this section we prove Theorem 3.1 from the main text, that claims that under some assumptions, for almost any initial configuration of particles, any normalized attention dynamics that we study (that is Post-LN, Pre-LN, Peri-LN, nGPT, Mix-LN and LN-Scaling) converges to a single cluster. First, let us outline the core of the proof.
@@ D.1  Proof Outline for Token Synchronization in Pre-LN
Existence of a Limit Point First, we demonstrate that the token dynamics indeed converge to a limiting configuration. This step heavily depends on the system. Common approach leverages the Łojasiewicz inequality, as seen in Geshkovski et al., (2025). It can be adopted to our setting, as we will show later. Moreover, our proof extends the gradient case $Q^{\top}K=V$ to a more general case, extending the synchronization results by Geshkovski et al., (2025), Criscitiello et al., (2024), even in Post-LN case.
Local behavior at the limiting point. Second, we must prove that any such limit point corresponds to the synchronized state where all tokens are identical. The classical argument involves a local stability analysis around the system’s critical points. One can typically show that any non-synchronized critical points are unstable and that their basin of attraction has measure zero, making them insignificant as final states. A comprehensive linearization analysis can be found in Criscitiello et al., (2024) that, in particular, covers Post-LN dynamics with $d\geq 3$. Together with a recent proof of synchronization for $d=2$ Polyanskiy et al., (2025), the stability of Post-LN system Jacobian is well-studied. We also rely on this method, but first we need to resolve the fact that Pre-LN system is non-compact.
@@ D.2  Generalized gradient descent convergence
First, we need to refine an important result of Łojasiewicz on convergence of gradient descent, so that it fits our problem setting. We follow an approach similar to the one presented in Haraux, (2012).
@@
When we get an autonomous dynamical system on a compact manifold, and we study its convergence to a limiting point, we need to study the Jacobian at that limiting point. Specifically, a well-known stability argument that was already written down several times (see Criscitiello et al., (2024), (Geshkovski et al.,, 2025, Lemma A.1)), employs central manifold theorem to show that basin of attraction of unstable critical points has measure zero.
@@
where $J_{qq}=-I_{n}$ and $J_{\theta q}=0$ because at the critical point $\nabla_{\theta_{j}}E(\Theta^{*})=0$. Therefore, it is enough to show that $J_{\theta\theta}$ has a positive eigenvalue. Because of the gradient-like structure, $J_{\theta\theta}$ is the product of two matrices – $\mathrm{diag}(\frac{1}{f_{1}(\Theta^{*})g_{1}(\Theta^{*})},\ldots,\frac{1}{f_{n}(\Theta^{*})g_{n}(\Theta^{*})})$ and a symmetric Hessian of the energy function $E$. The Hessian itself is unstable, this is an established result due to Criscitiello et al., (2024) (for $d\geq 3$) and Polyanskiy et al., (2025) (for $d=2$) that together closed synchronization for Post-LN.
@@
It is well-known that the set of unstable critical points of a dynamical system on a compact has measure-zero basin of attraction (see for example (Geshkovski et al.,, 2025, Lemma A.1) for a proof outline). Thus, we obtain that the event
@@
  • LN-Scaling Time change from Step 1 with $\tau=2\sqrt{t+1}$ reduces LN-Scaling to Post-LN.

### v2

@@ Appendix A Symmetric initialization
Post-LN. The ODE governing the evolution of the cosine similarity $\gamma(t)$ was already derived in (mathpersp23, Theorem 6.8). It is given by
@@
sqrt-scaling. The ODE governing $\gamma(t)$ is given by
@@
  • for sqrt-scaling $s_{k}\equiv\sqrt{t}$.
@@
  • $\mathcal{V}=e^{-4\sqrt{t}(1+o(1))}$ for sqrt-scaling.
@@ Appendix D Final convergence
In this section we prove Theorem 3.1 from the main text, that claims that under some assumptions, for almost any initial configuration of particles, any normalized attention dynamics that we study (that is Post-LN, Pre-LN, Peri-LN, nGPT, Mix-LN and sqrt-scaling) converges to a single cluster. First, let us outline the core of the proof.
@@ D.1  Proof Outline for Token Synchronization in Pre-LN
Existence of a Limit Point First, we demonstrate that the token dynamics indeed converge to a limiting configuration. This step heavily depends on the system. Common approach leverages the Łojasiewicz inequality, as seen in mathpersp23. It can be adopted to our setting, as we will show later. Moreover, our proof extends the gradient case $Q^{\top}K=V$ to a more general case, extending the synchronization results by mathpersp23, boumal, even in Post-LN case.
Local behavior at the limiting point. Second, we must prove that any such limit point corresponds to the synchronized state where all tokens are identical. The classical argument involves a local stability analysis around the system’s critical points. One can typically show that any non-synchronized critical points are unstable and that their basin of attraction has measure zero, making them insignificant as final states. A comprehensive linearization analysis can be found in boumal that, in particular, covers Post-LN dynamics with $d\geq 3$. Together with a recent proof of synchronization for $d=2$ yao2025circle, the stability of Post-LN system Jacobian is well-studied. We also rely on this method, but first we need to resolve the fact that Pre-LN system is non-compact.
@@ D.2  Generalized gradient descent convergence
First, we need to refine an important result of Łojasiewicz on convergence of gradient descent, so that it fits our problem setting. We follow an approach similar to the one presented in Haraux12.
@@
When we get an autonomous dynamical system on a compact manifold, and we study its convergence to a limiting point, we need to study the Jacobian at that limiting point. Specifically, a well-known stability argument that was already written down several times (see boumal, (mathpersp23, Lemma A.1)), employs central manifold theorem to show that basin of attraction of unstable critical points has measure zero.
@@
where $J_{qq}=-I_{n}$ and $J_{\theta q}=0$ because at the critical point $\nabla_{\theta_{j}}E(\Theta^{*})=0$. Therefore, it is enough to show that $J_{\theta\theta}$ has a positive eigenvalue. Because of the gradient-like structure, $J_{\theta\theta}$ is the product of two matrices – $\mathrm{diag}(\frac{1}{f_{1}(\Theta^{*})g_{1}(\Theta^{*})},\ldots,\frac{1}{f_{n}(\Theta^{*})g_{n}(\Theta^{*})})$ and a symmetric Hessian of the energy function $E$. The Hessian itself is unstable, this is an established result due to boumal (for $d\geq 3$) and yao2025circle (for $d=2$) that together closed synchronization for Post-LN.
@@
It is well-known that the set of unstable critical points of a dynamical system on a compact has measure-zero basin of attraction (see for example (mathpersp23, Lemma A.1) for a proof outline). Thus, we obtain that the event
@@
  • sqrt-scaling Time change from Step 1 with $\tau=2\sqrt{t+1}$ reduces sqrt-scaling to Post-LN.